If Levine, Inc., has an ROA of 7.8 percent and a payout ratio of 25 percent, what is its internal growth rate?  (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

Internal growth rate %
Answer #1

Retention ratio=1-payout ratio

=(1-0.25)=0.75

Internal growth rate=(ROA*Retention ratio)/[1-(ROA*Retention ratio)]

=(0.078*0.75)/[1-(0.078*0.75)]

=6.21%(Approx).